OpenStudy (anonymous):

The number of solutions of the equation 3p + 4q = 70, where p and q are positive integers and p > q, is

OpenStudy (lgbasallote):

hmm...this is linear isnt it?

OpenStudy (anonymous):

yup

OpenStudy (lgbasallote):

do you mean the possible values of p and q?

OpenStudy (anonymous):

yaa

OpenStudy (lgbasallote):

i would love to say there are infinite solutions because there are no dmain restrictions but i cant be sure

OpenStudy (zzr0ck3r):

I think either none, infinite or 1.

OpenStudy (anonymous):

@igbasallote how did u do that... any explanation

OpenStudy (lgbasallote):

lol didnt i just explain it =))

OpenStudy (anonymous):

\[3p + 4q = 70\] \[p = \frac{70 - 4q}{3}\] Now put the values of q like 1, 2, 3 etc and find corresponding q values.. p and q values must be integer.. First one I do for you: Put q = 1, \[p = \frac{70-4}{3} = \frac{66}{3} = 22\] So, one solution is : \[(p, q) = (22, 1)\] Likewise find others.. In total you will have three solutions..

OpenStudy (zzr0ck3r):

ahh nice work

OpenStudy (anonymous):

wow thanxx :)

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

two unknows and 1 equation..solution does make a sense here